OpenStudy (anonymous):

Solve (10^2x)^2 = 3^(x+1)

OpenStudy (xapproachesinfinity):

hmm is that \[(10^{2x})^2=3^{x+1}\]

OpenStudy (anonymous):

yass

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

ok so \[100^{x^2}=3^{x+1}\] we use log base 3 or loge base 100 \[x^2=\log_{100}(3^{x+1})\]

OpenStudy (xapproachesinfinity):

we then use change of base \[x^2=\frac{\log_3(3^{x+1})}{\log_3(100)}\]

OpenStudy (xapproachesinfinity):

can you continue>>>

OpenStudy (anonymous):

i dont get how you got log base 100?

OpenStudy (xapproachesinfinity):

i just applied log base 100 to get rid of 100 to power x^2 in the left

OpenStudy (xapproachesinfinity):

after i used log property. change of base

OpenStudy (anonymous):

ok so now what do you do?

OpenStudy (campbell_st):

ok... so start by using the index law for power of a power which means multiplying the powers so you get \[10^{4x} = 3^{x + 1}\] now take the base 10 log of both sides \[\log_{10}(10^{4x} )= \log_{10}(3^{x + 1})\] next apply the log laws for powers... \[4x \log_{10}(10) = (x + 1)\log_{10}(3)\] does that make sense so far..?

OpenStudy (campbell_st):

and in the next set you will need to know \[\log_{10}(10) = 1\]

OpenStudy (anonymous):

yass the first part makes sense, but i got lost on what you just put

OpenStudy (anonymous):

oh wait nevermind i got it

OpenStudy (anonymous):

so now we have 4x= x+1(log3)

OpenStudy (campbell_st):

ok... when you take the log of the base... its equal to 1.... \[10 = 10^1~~so~~~\log_{10}(10^1) = 1\] remember, using logs is about finding the power that when applied to the base gives the number needed... as a further example \[\log_{10}(100) = \log_{10}(10^2) = 2\] the base is 10, the power is 2 so 10^2 = 100 the starting number

OpenStudy (campbell_st):

so if you make the substitution \[\log_{10}(10) = 1\] then you get \[4x \times 1 = (x + 1)\log_{10}(3)\] distribute the right hand side \[4x = xlog_{10}(3) + \log_{10}(3)\] is that ok...?

OpenStudy (anonymous):

yeah i got that

OpenStudy (campbell_st):

ok... now collect the like terms \[4x - x \log_{10}(3) = \log_{10}(3)\] and factorise \[x(4 - \log_{10}(3)) = \log_{10}(3)\] does that make sense... and can you solve for x, from here...?

OpenStudy (anonymous):

yeah i get it thank yo so much, you would divide next right?

OpenStudy (campbell_st):

that's correct, by everything in the brackets on the left

OpenStudy (anonymous):

ok i got it thnx